OpenStudy (anonymous):

Im Confused What is y-13=4(x-2)

OpenStudy (anonymous):

y-13 = 4x - 8 y = 4x - 8 + 13 y = 4x + 5 An equation of a line..

OpenStudy (anonymous):

thank you